One number is 10 less than another. If 9 times the smaller number minus 2 times the larger number is 36, find the two numbers.
Hint: You must find both numbers for full credit.
Answer:
First number:
Second number:

Answer by ankor@dixie-net.com(22740)

:
let another number = x
:
"One number is 10 less than another".
One number = (x-10)
:
"If 9 times the smaller number minus 2 times the larger number is 36,"
(x - 10) = smaller number and x = the larger number
9(x-10) - 2x = 36
Do some basic algebra, solve for x
9x - 90 - 2x = 36
9x - 2x = 36 + 90
7x = 126
x = 126/7
x = 18,
:
"Find the two numbers"
Numbers would be 18 and 18-10 = 8:
:
Check using "9 times the smaller number minus 2 times the larger number is 36,"
9(8) - 2(18) =
72 - 36 = 36; proves our solution